### v3

# Fabien Le Floc’h

Abstract
We present closed analytical approximations for the pricing of Asian basket spread options under the Black-Scholes model. The formulae are obtained by using a stochastic Taylor expansion around a log-normal proxy model and are found to be highly accurate for Asian and spread options in practice. Unlike other approaches, they do not require any numerical integration or root solving.
keywords:
Stochastic expansion, Asian option, basket option, spread option, Black-Scholes, arithmetic averaging.
1 Introduction
The buyer of a vanilla basket spread option of strike $K$ and maturity $T$ on underlying assets $(S_{i})_{i=1,...,m}$ receives at maturity $T$ the payoff
(1) $V(T)=\max\left(\sum_{i=1}^{m}w_{i}S_{i}(T)-K,0\right)$
for a call and $\max\left(K-\sum_{i=1}^{m}w_{i}S_{i}(T),0\right)$ for a put. While for a vanilla basket option the weights $w_{i}$ are positive and the strike price obeys $K>0$, for a spread option the weights $w_{i}$ are negative for $i=1,...,m_{p}-1$, positive for $i=m_{p},...,n$ and are usually chosen such that $\sum_{i=1}^{m_{p}-1}w_{i}=-1$ and $\sum_{i=m_{p}}^{n}w_{i}=1$. The strike price $K$ may be positive or negative. Those are not so common in the domain of foreign exchange or equity derivatives (with $m>2$). The most popular example, the so-called crack spread stems from commodity derivatives. The portfolio of the 3:2:1 crack spread reads $3S_{3}-2S_{2}-S_{1}$, where $S_{3},S_{2}$ and $S_{1}$ are respectively the prices of crude oil, gasoline and heating oil. Other commodities such as soybean or cattle may also involve spreads on three different future contracts.
The buyer of an Asian call spread option of strike $K$ and averaging dates $(t_{j})_{j=1,...,n}$ on an asset $S$ receives
(2) $V(T)=\max\left(\sum_{j=1}^{n}w_{j}S(t_{j})-K,0\right)\,.$
Those are sometimes traded in equity and foreign exchange spaces. Under the Black-Scholes model with a term-structure of interest rates, dividend yields and volatilities, Asian spread options with discrete arithmetic averaging can be reduced to an equivalent vanilla basket spread option with a specific correlation structure.
Asian basket spread options combine both features: each underlying asset is averaged, typically over the same schedule and with the same weights, the payoff at maturity reads
(3) $V(T)=\max\left(\sum_{i=1}^{m}\sum_{j=1}^{n}w^{B}_{i}w^{A}_{j}S_{i}(t_{j})-K,0\right)\,,$
where $\bm{w}^{A}$ are positive Asianing weights and typically sum to 1, while $\bm{w}^{B}$ are basket weights of mixed sign. Asian basket spread options are mostly traded by commodity traders.
In the context of the Black-Scholes model, an early and popular approximation for prices of vanilla spread options (for $m=2$), is due to Kirk and Aron (1995). There exists however a simple semi-analytical exact formula based on a one-dimensional integral (Pearson, 1999), which should be preferred nowadays. As the number of assets $m$ (or the number of Asianing dates) increases, the integral formulation becomes less practical. Castellacci and Siclari (2003) combined the approximation of the arithmetic mean by a geometric mean (originally used by Vorst (1992) for vanilla Asian options) with the exchange option formula of Margrabe (1978) to obtain a simple approximation for the prices of Asian basket spread options. Deelstra et al. (2010) found approximations based on comonotonic bounds, which necessitate a one dimensional integration but compare favorably to the three moment matching shifted lognormal approximation of Borovkova et al. (2007). More recently, Pellegrino (2016) derived more accurate approximations, which may be seen as an extension of Bjerksund and Stensland (2014).
Here, we consider a Black-Scholes setting with time-dependent parameters, and propose a stochastic Taylor expansion using Margrabe exchange option formula on the geometric averages as a proxy, following the ideas of Etoré and Gobet (2012), who applied their expansion to price vanilla options with discrete dividends under the piecewise-lognormal model. It may be seen as a generalization of our previous work regarding vanilla basket and Asian options in (Le Floc’h, 2024). The resulting formula does not involve any numerical integration or root solving, thus offering computational speed and facilitating hedges calculations (for example using the option greeks). We start by deriving the proxy, the first-, second- and third-order expansions and then evaluate their accuracy on a sample of previously published cases. This allows for a comparison with other accurate approximations.
2 Reductions to vanilla basket spread options
2.1 From Asian to basket
Let $F(0,t)$ be the forward price of the underlying asset from valuation time to the date $t$. In the Black-Scholes model with time-dependent parameters we have
$F(0,t)=S(0)e^{\int_{0}^{t}r(s)-q(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\,,$
with $r,q$ being the instantateneous interest rate and dividend yield and $S(t)$ the underlying asset spot at time $t$. Under the risk-neutral measure $\mathbb{Q}$, the forward price follows
$\mathop{}\!\kern 0.0pt\mathrm{d}F(t,T)=F(t,T)\sigma(t)\mathop{}\!\kern 0.0pt\mathrm{d}W(t)\,,$
where $W$ is a $\mathbb{Q}$-Brownian motion.
The buyer of an arithmetic Asian option of strike $K$ and maturity $T$ on an underlying $S$ with observation dates $(t_{i})_{i=1,...,n}$ with $t_{n}=T$, receives at maturity $\max\left(\sum_{i=1}^{n}w_{i}S(t_{i})-K\right)$ for a call and $\max\left(K-\sum_{i=1}^{n}w_{i}S(t_{i})\right)$ for a put. It may be expressed as a basket option of strike $K$ and maturity $T$ on $n$ underlying assets $(S_{i})_{i=1,...,n}$. The covariance between the assets must match the covariance of $S$ between different observation times, this means [^1] We use the Itô isometry to obtain Equations 4 and 6.
(4) $\rho_{i,j}\sqrt{v_{i}v_{j}}=v_{i\wedge j}\,,$
where $\rho_{i,j}$ denotes the correlation between the assets $S_{i}$ and $S_{j}$, $v_{i}=\int_{0}^{t_{i}}\sigma^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$ being the total variance up to time $t_{i}$ for the asset $S$. Each underlying asset forward to $T$ must match the forwards of the single asset to $t_{i}$, meaning
(5) $F_{i}(0,T)=\mathbb{E}[S_{i}(T)]=\mathbb{E}[S(t_{i})]=F(0,t_{i})=S(0)e^{\int_{0}^{t_{i}}(r(s)-q(s))\mathop{}\!\kern 0.0pt\mathrm{d}s}\,.$
Thus, the arithmetic Asian option on $S$ with observations $t_{i}$ corresponds to a vanilla basket option on $S_{i}$ with forwards to maturity $F_{i}(0,T)$, total variances $v_{i}$ and correlations $\rho_{i,j}$ for $i,j=1,...,n$.
2.2 Asian baskets
This may be extended to Asian baskets with payoff at maturity
$\max\left(\sum_{i=1}^{n}w_{i}\sum_{j=1}^{m}\mu_{j}S_{j}(t_{i})-K,0\right)\,,$
where $(\mu_{j})_{j=1...m}$ are basket weights and $S_{j}$ are the different assets with correlation $\rho_{j,l}$. We may define $n\times m$ assets $\hat{S}_{i+n(j-1)}$ with forward $F_{j}(t_{i})$, total variance $\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$, correlation
(6) $\hat{\rho}_{i+n(j-1),k+n(l-1)}=\rho_{j,l}\frac{\int_{0}^{t_{i}\wedge t_{k}}\sigma_{j}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\int_{0}^{t_{k}}\sigma_{l}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}}\,,$
for $(i,k)\in\{1,...,n\}^{2}$, $(j,l)\in\left\{1,...,m\right\}^{2}$. The Asian basket option is equivalent to a vanilla basket option on $\hat{S}_{i+n(j-1)}$ with weights $w_{i}\mu_{j}$.
2.3 Strike
The strike $K$ may be included in the spread sum. In the context of an Asian, it is equivalent to an asian observation at $t=0$ with weight $-K/S(0)$. In the context of a basket, it is equivalent to an additional asset of spot price $K$ with zero volatility and drift, weight $-1$, and correlation zero towards the other assets.
3 Expansions
3.1 The proxy
An early technique, first explored by Vorst (1992) to estimate the price of Asian options is to approximate the arithmetic average by the geometric average, because the geometric average of geometric Brownian motions is another geometric Brownian motion and can thus be priced with the usual Black-Scholes formula. This was extended to the more general cases of vanilla basket options by Gentle (1993).
Under the risk-neutral measure $\mathbb{Q}$, the forward price for each asset follows
$\mathop{}\!\kern 0.0pt\mathrm{d}F_{i}(t,T)=F_{i}(t,T)\sigma_{i}(t)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(t)\,,$
where $W_{i}$ is a $\mathbb{Q}$-Brownian motion, and we have $\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}\mathop{}\!\kern 0.0pt\mathrm{d}W_{j}=\rho_{i,j}\mathop{}\!\kern 0.0pt\mathrm{d}t$. The payoff at maturity of the basket spread option where we allow for a multiplicative strike $\kappa\in\mathbb{R}$ reads
$V(T)$ $=\left[\eta\left(\sum_{i=m_{p}}^{m}w_{i}S_{i}(T)+\kappa\sum_{i=1}^{m_{p}-1}w_{i}S_{i}(T)\right)\right]^{+}$
$=A_{p}\left[\eta\left(\sum_{i=m_{p}}^{m}\tilde{a}_{i}S_{i}^{\star}(T)-{\kappa}^{\star}\sum_{i=1}^{m_{p}-1}|\tilde{a}_{i}|S_{i}^{\star}(T)\right)\right]^{+}$
with $\eta=\pm 1$ for a call (respectively a put) option and
$A_{n}$ $=-\sum_{i=1}^{m_{p}-1}w_{i}F_{i}(0,T)\,,\quad A_{p}=\sum_{i=m_{p}}^{m}w_{i}F_{i}(0,T)\,,\quad\kappa^{\star}=\frac{A_{n}}{A_{p}}\kappa\,,$
$\tilde{a}_{i}$ $=\frac{w_{i}F_{i}(0,T)}{A_{p}(T)}\quad\textmd{ for }i\geq m_{p}\quad\textmd{ and }\quad\tilde{a}_{i}=\frac{w_{i}F_{i}(0,T)}{A_{n}(T)}\quad\textmd{ for }i<m_{p}\,,$
$S_{i}^{\star}(t)$ $=\frac{S_{i}(t)}{F_{i}(0,t)}=e^{-\frac{1}{2}\int_{0}^{t}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}\,.$
In particular, we have $\tilde{a}_{i}<0$ for $i<m_{p}$, and $A_{n}>0$. We approximate the arithmetic average by the geometric average leading to
(7) $V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\prod_{i=m_{p}}^{m}S_{i}^{\star}(T)^{a_{i}}-{\kappa}^{\star}\prod_{i=1}^{m_{p}-1}S_{i}^{\star}(T)^{|a_{i}|}\right)\right)^{+}\right]$
with $a_{i}=\tilde{a}_{i}$ as a specific case.
Let
$G(T,1,n)=\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}=e^{\sum_{i=1}^{n}-\frac{1}{2}a_{i}\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{T}a_{i}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}$
We have
(8) $m(T,1,n)=\mathbb{E}\left[\ln G(T)\right]$ $=-\frac{1}{2}\sum_{i=1}^{n}a_{i}v_{i}\,,$
(9) $\tilde{\nu}^{2}(T,1,n)=\mathrm{Var}\left[\ln G(T)\right]$ $=\sum_{i,j=1}^{n}a_{i}a_{j}\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
with $v_{i}=\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$. We will also define the total variances of the positive and negative parts
(10) $\tilde{\nu}_{n}^{2}(T)=\tilde{\nu}(T,1,m_{p}-1)\,,\quad\tilde{\nu}_{p}^{2}=\tilde{\nu}(T,m_{p},m)\,,\quad\tilde{\nu}^{2}(T)=\tilde{\nu}(T,1,m)\,.$
Instead of adjusting the strike to preserve the expectation of the term inside the max function as in (Vorst, 1992), we adjust each geometric average by a factor (respectively $\alpha$ and $\beta$) to match the first moment of each corresponding arithmetic average:
(11) $V(0)\approx B(0,T)A_{p}\mathbb{E}\left[\left(\eta\left(\alpha\prod_{i=m_{p}}^{m}S_{i}^{\star}(T)^{a_{i}}-{\kappa}^{\star}\beta\prod_{i=1}^{m_{p}-1}S_{i}^{\star}(T)^{|a_{i}|}\right)\right)^{+}\right]\,,$
with $\alpha=\frac{1}{\mathbb{E}[G(T,m_{p},m)]}$, $\beta=\mathbb{E}[G(T,1,m_{p}-1)]$ and $B(0,T)$ is the discount factor to the payment date.
Levy (1992) matches not only the mean but also the variance of the arithmetic average by a lognormal distribution. We have
(12) $\nu_{A_{p}}^{2}$ $=\ln\left(\mathrm{Var}\left[A_{p}(T)\right]\right)=\ln\left(\sum_{i,j=m_{p}}^{m}a_{i}a_{j}e^{\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\right)\,,$
(13) $\nu_{A_{n}}^{2}$ $=\ln\left(\mathrm{Var}\left[A_{n}(T)\right]\right)=\ln\left(\sum_{i,j=1}^{m_{p}-1}a_{i}a_{j}e^{\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\right)\,.$
We may adjust the $a_{i}$ used in the geometric mean such that the variance of the geometric mean matches exactly the one of the arithmetic mean by using
(14) $a_{i}=\tilde{a}_{i}\frac{\nu_{A_{p}(T)}}{\tilde{\nu}_{p}(T)}\quad\textmd{ for }i\geq m_{p}\,,\quad\textmd{ and }\quad a_{i}=\tilde{a}_{i}\frac{\nu_{A_{n}}}{\tilde{\nu}_{n}(T)}\quad\textmd{ for }i<m_{p}\,.$
In Identity 1, we apply a change of measure to reduce Equation 11 to a regular vanilla option pricing problem as in (Margrabe, 1978). This leads to
(15) $V(0)\approx A_{p}\cdot\mathrm{Black}\left(1,\kappa^{\star},{\nu}^{2},T\right)$
where $\mathrm{Black}(F,K,v,T)$ denotes the Black-76 formula applied to the forward $F$, strike $K$ and total variance $v$ to maturity $T$ (Equation 21) and $\nu=\tilde{\nu}$ for the geometric approximation, $\nu$ computed from $a_{i}$ for the Levy approximation.
We will adopt the Vorst geometric formula with adjusted forward as proxy for our stochastic Taylor expansion. Instead of our normalization factors $\alpha$ and $\beta$, Krekel (2003) employs an additive shift to adjust the expectation of the whole spread, as does Vorst (1992) for regular Asian options. It would lead to slightly more complicated equations and exhibit a worse accuracy.
In the context of standard Asian options (no spread), the Vorst Levy formula corresponds exactly to the Levy approximation, but defines a stochastic process based on the individual underlying asset processes. This will allow to define stochastic expansions based on this alternative proxy as well.
Both approximations have a major advantage compared to a shifted lognormal moment matching approach, such as in (Borovkova et al., 2007): each sum of lognormal distributions is approximated by a distribution defined on the positive real line. In contrast, for example in the case of a basket with positive weights, the shifted lognormal moment matching may lead to a positive shift, which will allow for negative values of the basket price, which is not realistic. With weights of mixed signs, the shifted lognormal approximation requires special handling to capture negative skews, as described in (Borovkova et al., 2007).
3.2 First order expansion
Let $h(x)=\max(\eta x,0)$, we apply a Taylor expansion of order-1 on $h$:
(16) $\mathbb{E}\left[h\left(\sum_{i=m_{p}}^{m}\tilde{a}_{i}S_{i}^{\star}(T)-\kappa^{\star}\sum_{i=1}^{m_{p}-1}|\tilde{a}_{i}|S_{i}^{\star}(T)\right)\right]=\mathbb{E}\left[h\left(G_{p}^{\star}(T)-\kappa^{\star}G_{n}^{\star}(T)\right)\right]\\ +\mathbb{E}\left[\left(\sum_{i=m_{p}}^{m}\tilde{a}_{i}S_{i}^{\star}(T)-\kappa^{\star}\sum_{i=1}^{m_{p}-1}|\tilde{a}_{i}|S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)h^{\prime}\left(G_{p}^{\star}(T)-\kappa^{\star}G_{n}^{\star}(T)\right)\right]\\ +\mathrm{Error}_{2}(h)$
where $\mathrm{Error}_{2}(h)\leq_{c}2\left(||\sum_{i=m_{p}}^{m}\tilde{a}_{i}S_{i}^{\star}(T)-\kappa^{\star}\sum_{i=1}^{m_{p}-1}|\tilde{a}_{i}|S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)||_{3}\right)^{2}$.
Let $\tilde{a}_{i}^{\star}=\tilde{a}_{i}$ for $i\geq m_{p}$ and $\tilde{a}_{i}^{\star}=\kappa^{\star}\tilde{a}_{i}$ for $i<m_{p}$, we may rewrite the second term using a derivative with regards to $\kappa^{\star}$, and interchange derivation and expectation (see Etoré and Gobet (2012)) to obtain
(17) $\mathbb{E}\left[\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)h^{\prime}\left(G_{p}^{\star}(T)-\kappa^{\star}G_{n}^{\star}(T)\right)\right]\\ =-\frac{\partial}{\partial\kappa^{\star}}\mathbb{E}\left[\frac{\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)}{G_{n}^{\star}(T)}h\left(G_{p}^{\star}(T)-\kappa^{\star}G_{n}^{\star}(T)\right)\right]$
We use Identities 1, 2 and 3 to obtain the first order expansion
(18) $V_{\textmd{VG1}}=A_{p}\cdot\mathrm{Black}(1,\kappa^{\star},\nu^{2},T)\\ +A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\kappa^{\star}A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(1,\kappa^{\star},\nu^{2},T\right)\\ -A_{p}\sum_{i=1}^{n}\tilde{a}_{i}^{\star}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\,.$
with $\bar{v}_{i}$ defined in Equation 31. The relevant derivatives of the Black-76 formula are given in Appendix B. The formula with the Vorst Levy proxy is exactly the same, but uses different $\nu$ and $a_{i}$.
Interestingly, while our derivation is different, the first order expansion on the geometric average ”VG1” is the same as the extended Bjersksund Stensland approximation of Pellegrino (2016).
3.3 Second order expansion
The second-order Taylor expansion involves the additional term
$\frac{1}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathbb{E}\left[\frac{\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)^{2}}{{G_{n}^{\star}}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]$
We decompose it into six terms:
$\frac{\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)^{2}}{{G_{n}^{\star}}(T)}=G_{p}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}+{\kappa^{\star}}^{2}G_{n}^{\star}(T)\\ -2\kappa^{\star}G_{p}^{\star}(T)-2\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)\right)\frac{G_{p}^{\star}}{G_{n}^{\star}}+2\kappa^{\star}\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)\right)+\sum_{i=1}^{m}\sum_{j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\frac{S_{i}^{\star}(T)S_{J}^{\star}(T)}{G_{n}^{\star}(T)}\,.$
We then use Identities 1 to 6 to find our formula for the second order expansion:
(19) $V_{\textmd{VG2}}=V_{\textmd{VG1}}+\frac{A_{p}e^{\nu^{2}(T)}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)+\frac{\kappa^{2}A_{p}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\kappa^{\star}A_{p}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\kappa^{\star}{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$
with $\bar{v}_{i}$, $v_{i,j}$ defined in Equation 31 and $\bar{v}_{i}^{-}$ defined in Equation 36.
3.4 Third order expansion
The third order expansion adds the term
$-\frac{1}{6}\frac{\partial^{3}}{\partial{\kappa^{\star}}^{3}}\mathbb{E}\left[\frac{\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)^{3}}{{G_{n}^{\star}(T)}^{2}}h\left(\alpha\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]$
We decompose it into the following terms:
$\frac{\left(\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}(T)-G_{p}^{\star}(T)+\kappa^{\star}G_{n}^{\star}(T)\right)^{3}}{{G_{n}^{\star}(T)}^{2}}=-\frac{{G_{p}^{\star}}^{3}}{{G_{n}^{\star}}^{2}}+3\kappa^{\star}\frac{{G_{p}^{\star}}^{2}}{{G_{n}^{\star}}}-3{\kappa^{\star}}^{2}G_{p}^{\star}+{\kappa^{\star}}^{3}{G_{n}^{\star}}\\ +3\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{S_{i}^{\star}{G_{p}^{\star}}^{2}(T)}{{G_{n}^{\star}}^{2}}+3{\kappa^{\star}}^{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}S_{i}^{\star}-6\kappa^{\star}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{S_{i}^{\star}{G_{p}^{\star}}(T)}{{G_{n}^{\star}}}\\ -3\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\frac{S_{i}^{\star}S_{j}^{\star}G_{p}^{\star}(T)}{{G_{n}^{\star}}^{2}}+3\kappa^{\star}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\frac{S_{i}^{\star}S_{j}^{\star}}{{G_{n}^{\star}}}+\sum_{i,j,k=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\tilde{a}_{k}^{\star}\frac{S_{i}^{\star}S_{j}^{\star}S_{k}^{\star}(T)}{{G_{n}^{\star}}^{2}}$
We use Identities 1 to 10 to obtain the third-order formula
(20) $V_{\textmd{VG3}}=V_{\textmd{VG2}}\\ +\frac{A_{p}e^{3\nu^{2}(T)}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{3\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}\kappa^{\star}e^{\nu^{2}}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}{\kappa^{\star}}^{2}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}{\kappa^{\star}}^{3}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\frac{A_{p}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu^{2}+2\nu_{n}^{2}-2\nu_{pn}^{2}+2\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}{\kappa^{\star}}^{2}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +A_{p}{\kappa^{\star}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{3\nu_{n}^{2}-2\nu_{pn}^{2}+\bar{v}_{i}+\bar{v}_{j}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\nu^{2}-2\nu_{pn}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}\kappa^{\star}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}}{6}\sum_{i,j,k=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\tilde{a}_{k}^{\star}e^{{v}_{i,j}+{v}_{k,i}+{v}_{k,j}+2\bar{v}_{i}^{-}+2\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+3\nu_{n}^{2}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{k}-3\nu_{np}^{2}+3\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$
with $\bar{v}_{i}$, $v_{i,j}$ defined in Equation 31 and $\bar{v}_{i}^{-}$ defined in Equation 36.
3.5 Symmetries
We may exploit the symmetries in the double and triple sums to reduce the number terms being summed and speed up the corresponding algorithm. In particular,
$\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)=\sum_{i=1}^{n}\tilde{a}_{i}^{2}e^{v_{i,i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ +2\sum_{i=1}^{n}\sum_{j=1}^{i-1}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)\,.$
and
$\sum_{i,j,l=1}^{n}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)=\\ 6\sum_{i=1}^{n}\sum_{j=1}^{i-1}\sum_{l=1}^{j-1}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\\ +3\sum_{i=1}^{n}\sum_{j=1}^{i-1}\tilde{a}_{i}\tilde{a}_{j}^{2}e^{2v_{i,j}+v_{j,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+2\bar{v}_{j}},K^{\star},\nu^{2},T\right)\\ +3\sum_{i=1}^{n}\sum_{l=1}^{i-1}\tilde{a}_{i}^{2}\tilde{a}_{l}e^{2v_{i,l}+v_{i,i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{2\bar{v}_{i}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\\ +\sum_{i=1}^{n}\tilde{a}_{i}^{3}e^{3v_{i,i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{3\bar{v}_{i}},K^{\star},\nu^{2},T\right)\,.$
4 Numerical examples
For regular Asian and basket options (i.e., with positive weights only), the approximations are equivalent to the ones of Le Floc’h (2024). We will thus consider only cases of spread options, (i.e., with some negative weights).
4.1 Spread options
We start with the example presented in (Deelstra et al., 2010, Table 1). It consists in pricing a call spread option of maturity 1 year paying $(S_{2}-S_{1}-K)^{+}$, using a risk-free interest rate $r=5\%$ on two assets $S_{1}$ and $S_{2}$ with $S_{1}(0)=200$, $S_{2}(0)=100$ and $\sigma_{1}=60\%,\sigma_{2}=60\%$ with correlation $\rho_{1,2}=28\%$ for a range of strikes $K$. Our Monte-Carlo prices differs slightly from the ones presented in Deelstra et al. (2010) as we employed a Monte-Carlo simulation with a digitally scrambled Sobol sequence (L’Ecuyer and Lemieux, 2002) on 64 millions paths. Compared to the Monte-Carlo simulation of Deelstra et al. (2010) using 100 millions of paths, our prices have a better accuracy, which is important since they are used to compute the root mean square errors (RMSE) and maximum absolute errors (MAE) of the various approximations. On this example, the basket skewness $\mu_{3}$ is negative, where
$\mu_{3}=\frac{\mathbb{E}\left[\left(\sum_{i=1}^{m}w_{i}S_{i}(T)-\mathbb{E}\left[\sum_{i=1}^{m}w_{i}S_{i}(T)\right]\right)^{3}\right]}{\left(\mathbb{E}\left[\left(\sum_{i=1}^{m}w_{i}S_{i}(T)\right)^{2}\right]-\mathbb{E}\left[\sum_{i=1}^{m}w_{i}S_{i}(T)\right]^{2}\right)^{\frac{3}{2}}}\,.$
As per Table 1, the proxy prices (VG0) are already quite accurate, much better than the shifted lognormal (SLN) approximation of Borovkova et al. (2007). The first order (VG1) does not improve the accuracy, while the second order (VG2) is an order of magnitude more accurate. The third order (VG3) allows to gain another order of magnitude in accuracy. The most accurate is the improved comonotonic upper bound (ICUB) approximation of Deelstra et al. (2010).
Table 1: Call spread option of maturity 1 year paying $(S_{2}-S_{1}-K)^{+}$, on two assets $S_{1}$ and $S_{2}$ with $S_{1}(0)=200$, $S_{2}(0)=100$ and $\sigma_{1}=60\%,\sigma_{2}=60\%$ with correlation $\rho_{1,2}=28\%$. This example has negative basket skewness
$K$ VG0 VG1 VG2 VG3 SLN ICUB MC
-70 28.9222 28.9251 29.0906 29.0853 31.0619 29.0854 29.0854
-80 33.4520 33.4313 33.6233 33.6148 35.8096 33.6150 33.6150
-90 38.3631 38.3258 38.5397 38.5276 40.8772 38.5281 38.5281
-100 43.6358 43.5882 43.8192 43.8031 46.2500 43.8043 43.8043
-110 49.2485 49.1961 49.4394 49.4194 51.9127 49.4212 49.4212
-120 55.1792 55.1263 55.3770 55.3535 57.8497 55.3561 55.3561
-130 61.4056 61.3551 61.6093 61.5826 64.0453 61.5861 61.5861
RMSE 0.1701 0.2084 0.0158 0.0018 2.3512 0.0000 0.0000
MAE 0.1805 0.2310 0.0232 0.0034 2.4936 0.0000 0.0000
The second example originally from (Deelstra et al., 2010, Table 2) is a similar spread option but using $S_{1}(0)=40,S_{2}(0)=100$, $\sigma_{1}=17\%,\sigma_{2}=40\%$ and $\rho_{1,2}=12\%$. The smaller vols results in a much more accurate Taylor expansion, and the second order expansion is the most accurate on this example (Table 2). The ICUB approximation is less accurate as in the context of a negative basket skewness.
Table 2: Call spread option of maturity 1 year paying $(S_{2}-S_{1}-K)^{+}$, on two assets $S_{1}$ and $S_{2}$ with $S_{1}(0)=40,S_{2}(0)=100$, $\sigma_{1}=17\%,\sigma_{2}=40\%$ and $\rho_{1,2}=12\%$. This example has positive basket skewness.
$K$ VG0 VG1 VG2 VG3 SLN ICUB MC
45 24.5969 24.5971 24.5982 24.5982 24.6096 24.5975 24.5982
50 21.8235 21.8236 21.8247 21.8247 21.8441 21.8240 21.8247
55 19.3074 19.3074 19.3086 19.3086 19.3342 19.3079 19.3086
60 17.0380 17.0380 17.0391 17.0391 17.0692 17.0384 17.0391
65 15.0018 15.0017 15.0029 15.0029 15.0355 15.0022 15.0029
70 13.1831 13.1830 13.1842 13.1842 13.2178 13.1835 13.1842
75 11.5653 11.5652 11.5663 11.5663 11.5997 11.5656 11.5663
RMSE 0.0011 0.0011 0.0000 0.0000 0.0277 0.0007 0.0000
MAE 0.0013 0.0012 0.0000 0.0000 0.0336 0.0007 0.0000
4.2 Basket spread options
We now consider the example from (Deelstra et al., 2010, Table 4) of a basket spread option with payoff $(S_{3}-S_{2}-S_{1}-K)^{+}$ at maturity $T=1$, where $S_{3}(0)=100$, $S_{2}(0)=24$, $S_{1}(0)=46$, $\sigma_{3}=40\%$, $\sigma_{2}=22\%$, $\sigma_{1}=30\%$, $\rho_{3,2}=17\%$, $\rho_{1,3}=91\%$, $\rho_{2,1}=41\%$. On this first example (Table 3), the second-order and third-order reach beyond our Monte-Carlo reference prices accuracy.
Table 3: Basket spread option with payoff $(S_{3}-S_{2}-S_{1}-K)^{+}$ at maturity $T=1$, where $S_{3}(0)=100$, $S_{2}(0)=24$, $S_{1}(0)=46$, $\sigma_{3}=40\%$, $\sigma_{2}=22\%$, $\sigma_{1}=30\%$, $\rho_{3,2}=17\%$, $\rho_{1,3}=91\%$, $\rho_{2,1}=41\%$.
$K$ VG0 VG1 VG2 VG3 SLN HICUB MC
25.0 20.7593 20.7611 20.7646 20.7646 20.8073 20.7637 20.7645
30.0 17.6876 17.6884 17.6931 17.6931 17.7711 17.6921 17.6931
35.0 14.9531 14.9532 14.9590 14.9589 15.0630 14.9580 14.9591
40.0 12.5514 12.5512 12.5577 12.5576 12.6769 12.5566 12.5579
45.0 10.4677 10.4676 10.4746 10.4744 10.5983 10.4734 10.4747
50.0 8.6800 8.6799 8.6871 8.6870 8.8065 8.6859 8.6873
55.0 7.1612 7.1613 7.1684 7.1682 7.2767 7.1671 7.1685
RMSE 0.0829 0.0054 0.0001 0.0000 0.0455 0.0959 0.0000
MAE 0.1351 0.0059 0.0002 0.0000 0.0566 0.1625 0.0000
Next, we consider a similar example (Deelstra et al., 2010, Table 7), presented in Table 4 with higher volatilities: $S_{3}(0)=100$, $S_{2}(0)=63$, $S_{1}(0)=12$, $\sigma_{3}=21\%$, $\sigma_{2}=34\%$, $\sigma_{1}=63\%$, $\rho_{3,2}=87\%$, $\rho_{1,3}=30\%$, $\rho_{2,1}=43\%$.
Table 4: Basket spread option with payoff $(S_{3}-S_{2}-S_{1}-K)^{+}$ at maturity $T=1$, where $S_{3}(0)=100$, $S_{2}(0)=63$, $S_{1}(0)=12$, $\sigma_{3}=21\%$, $\sigma_{2}=34\%$, $\sigma_{1}=63\%$, $\rho_{3,2}=87\%$, $\rho_{1,3}=30\%$, $\rho_{2,1}=43\%$.
$K$ VG0 VG1 VG2 VG3 SLN HICUB MC
2.5 23.4535 23.5493 23.5998 23.5941 23.1681 23.5138 23.5938
10.0 17.0131 17.1596 17.2144 17.2040 16.8591 17.1373 17.2063
17.5 11.2434 11.3588 11.4175 11.4084 11.3394 11.3873 11.4112
25.0 6.5548 6.5418 6.6041 6.6006 6.9203 6.6584 6.6023
32.5 3.2670 3.1203 3.1849 3.1870 3.7629 3.3147 3.1877
40.0 1.3636 1.1852 1.2504 1.2527 1.7925 1.3853 1.2518
47.5 0.4758 0.3537 0.4072 0.4032 0.7369 0.4913 0.4024
RMSE 0.1263 0.0560 0.0050 0.0016 0.4041 0.0900 0.0000
MAE 0.1932 0.0675 0.0081 0.0028 0.5752 0.1335 0.0000
The second order expansion is more accurate than the hybrid ICUB by a factor of 10. The latter was found to be the most accurate in (Deelstra et al., 2010). The third order expansion further improves accuracy by a factor four.
4.3 Asian basket spread options
The option’s maturity is set equal to one year and the averaging is taken over the last 30 days, which is a common practice in energy markets.
The first example comes from (Deelstra et al., 2010, Table 8) and consists in an Asian basket spread option with payoff $\left(\sum_{i=1}^{n}w_{i}S_{2}(t_{i})-\sum_{i=1}^{n}w_{i}S_{1}(t_{i})-K\right)^{+}$ and $S_{2}(T)=100$, $S_{1}(T)=60$, $\sigma_{2}=33\%$, $\sigma_{1}=25\%$, $\rho_{1,2}=40\%$, $w_{i}=1/30$, $t_{i}=T-(i-1)/30$, $n=30$. In order to assess the true accuracy of the third-order expansion, we recalculated the reference Monte-Carlo prices with a higher accuracy than Deelstra et al. (2010), by using Brownian-Bridge variance reduction on a randomized Sobol sequence: the third order expansion reaches the limits of our quasi Monte-Carlo simulation accuracy with more than 32 millions paths (5).
Table 5: Asian basket spread option with payoff $\left(\sum_{i=1}^{n}w_{i}S_{2}(t_{i})-\sum_{i=1}^{n}w_{i}S_{1}(t_{i})-K\right)^{+}$ with $S_{2}(T)=100$, $S_{1}(T)=60$, $\sigma_{2}=33\%$, $\sigma_{1}=25\%$, $\rho_{1,2}=40\%$.
$K$ VG0 VG1 VG2 VG3 SLN HICUB MC
25.0 20.7593 20.7611 20.7646 20.7646 20.8073 20.7637 20.7646
30.0 17.6876 17.6884 17.6931 17.6931 17.7711 17.6921 17.6931
35.0 14.9531 14.9532 14.9590 14.9589 15.0630 14.9580 14.9589
40.0 12.5514 12.5512 12.5577 12.5576 12.6769 12.5566 12.5577
45.0 10.4677 10.4676 10.4746 10.4744 10.5983 10.4734 10.4745
50.0 8.6800 8.6799 8.6871 8.6870 8.8065 8.6859 8.6870
55.0 7.1612 7.1613 7.1684 7.1682 7.2767 7.1671 7.1683
RMSE 0.0063 0.0060 0.0001 0.0000 0.1031 0.0010 0.0000
MAE 0.0070 0.0070 0.0002 0.0000 0.1238 0.0012 0.0000
The next example corresponds to (Deelstra et al., 2010, Table 10) and is an Asian basket spread option with payoff
$V(T)=\left(\sum_{i=1}^{n}w_{i}S_{3}(t_{i})-\sum_{i=1}^{n}w_{i}S_{2}(t_{i})-\sum_{i=1}^{n}w_{i}S_{1}(t_{i})-K\right)^{+}\,,$
with $S_{3}(0)=100$, $S_{2}(0)=50$, $S_{1}(0)=25$, $\sigma_{3}=35$, $\sigma_{2}=30\%$, $\sigma_{1}=25\%$, $\rho_{2,3}=30\%$, $\rho_{1,3}=80\%$, $\rho_{1,2}=70\%$.
On this second example, the proxy is already as accurate as the hybrid ICUB approximation, and the second order expansion reaches the limits of the Monte-Carlo simulation with 300 millions of paths from Deelstra et al. (2010), while the third-oder reaches the limits of our quasi Monte-Carlo simulation (Table 6). This confirms the findings of Le Floc’h (2024), which noticed that stochastic Taylor expansions were particularly accurate for Asian options, as the averaging reduces the effective volatility.
Table 6: Asian basket spread option with payoff $\left(\sum_{i=1}^{n}w_{i}(S_{3}(t_{i})-S_{2}(t_{i})-S_{1}(t_{i}))-K\right)^{+}$.
$K$ VG0 VG1 VG2 VG3 SLN HICUB MC
10.0 20.5577 20.5991 20.6010 20.6010 20.7157 20.5521 20.6010
15.0 17.5270 17.5514 17.5546 17.5545 17.7346 17.5209 17.5546
20.0 14.8298 14.8364 14.8409 14.8409 15.0677 14.8236 14.8409
25.0 12.4617 12.4516 12.4575 12.4574 12.7097 12.4559 12.4574
30.0 10.4079 10.3836 10.3906 10.3906 10.6478 10.4030 10.3906
35.0 8.6463 8.6108 8.6187 8.6186 8.8635 8.6425 8.6186
40.0 7.1499 7.1063 7.1147 7.1145 7.3342 7.1472 7.1145
RMSE 0.0270 0.0059 0.0001 0.0000 0.2188 0.0283 0.0000
MAE 0.0434 0.0082 0.0002 0.0000 0.2572 0.0489 0.0000
4.4 Asian spread options
We consider the example from Krekel (2003) of equity Asian spread options with payoff $\left(\sum_{i=27}^{52}w_{i}S(t_{i})-K\sum_{j=1}^{26}w_{j}S(t_{j})\right)^{+}$ where $w_{i}=1/26$, $t_{i}=\frac{i}{52}$ for $i\leq 26$ and $t_{i}=4.5+\frac{i-26}{52}$ for $i>26$, spot price $S(0)=100$, interest rate $r=5\%$, dividend yield $q=1\%$, for a range of multiplicative strikes $K$. We consider the most challenging case, that is the seasoned option (at $t=0.25$) with high volatility $\sigma=70\%$. Thus, thirteen observations are known, and their fixings are set to the original forward prices $F(0,t_{j})$. In this example, the forward prices are all set from $t=0$ and not from $t=0.25$. We compute reference Monte-Carlo prices using Brownian-bridge variance reduction on a digitally scrambled Sobol sequence of 1 billion paths.
The proxy is already more accurate than their moment matching formula [^2] Their moment matching formula consists in a one-dimensional integration of two lognormal moment matched distributions, one for the averaging spot part, one for the averaging strike part. ”MM” , which is the most accurate method presented in (Krekel, 2003, Table 2). The second order expansion is around 40 times more accurate, and the third order reaches the limit of our Monte-Carlo simulation accuracy, with a maximum error below $0.002$ basis point of the spot price (Table 7).
Table 7: Prices of seasoned Asian spread options with payoff $\left(\sum_{i=27}^{52}w_{i}S(t_{i})-K\sum_{j=1}^{26}w_{j}S(t_{j})\right)^{+}$ for a range of multiplicative strikes $K$.
$K$ VG0 VG1 VG2 VG3 MM QMC
0.5 68.69262 68.68561 68.68818 68.68812 68.69909 68.68810
0.8 59.55470 59.54542 59.54909 59.54901 59.56331 59.54899
1.0 54.79849 54.78831 54.79251 54.79242 54.80797 54.79240
1.2 50.77850 50.76772 50.77233 50.77224 50.78856 50.77223
1.5 45.76550 45.75420 45.75928 45.75919 45.77606 45.75917
RMSE 0.00582 0.00402 0.00010 0.00002 0.01497 0.00000
MAE 0.00633 0.00497 0.00011 0.00002 0.01689 0.00000
Those observations also holds true for the newly issued Asian spread option case.
5 Conclusion
The second-order expansion is found to be the most accurate, by far, compared to previously published approximations. The third-order expansion goes further beyond, and reaches the limits of our reference quasi Monte-Carlo simulation accuracy for a much lower computational cost. This shows the power of stochastic Taylor expansions: they are very simple and yet lead to surprisingly accurate approximations. Their range of applicability is also great. So far, we have covered discrete cash dividends, Asian options, basket options (to some extent) and Asian basket spread options under the Black-Scholes model with a term-structure of rates and volatilities.
Further research may explore stochastic Taylor expansions with assets under stochastic volatility dynamics. The change of measure may still lead to explicit formulae for models with a known characteristic function.
Appendix A Black-Scholes identities
Let $\mathrm{Black}(F,K,v,T)$ denotes the Black-76 formula applied to the forward $F$, strike $K$ and total variance $v$ to maturity $T$:
(21) $\mathrm{Black}(F,K,v,T)=\eta B(0,T)\left[F\Phi(\eta d_{1})-K\Phi(\eta d_{2})\right]$
with $d_{1}=\frac{1}{\sqrt{v}}\left(\ln\frac{F}{K}+\frac{1}{2}v\right)$, $d_{2}=d_{1}-\sqrt{v}$.
We also define
(22) $h(x)=\max(\eta x,0)$
which constitutes the basis of our Taylor expansions.
Identity 1.
Let $G_{n}^{\star}=\frac{\beta}{G(T,1,m_{p}-1)}$ and $G_{p}^{\star}=\alpha G(T,m_{p},m)$, we have
(23) $\mathbb{E}\left[B(0,T){G_{n}}^{\star}(T)h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\mathrm{Black}\left(1,\kappa^{\star},{\nu}^{2},T\right)\,,$
where $\nu=\tilde{\nu}$ for the geometric approximation, $\nu$ computed from $a_{i}$ for the Levy approximation.
Proof.
We have $\mathbb{E}\left[G_{n}^{\star}(T)\right]=1$ and $G_{n}^{\star}$ may be thus expressed as
[…]
in probability, with $W_{G_{n}}$ being a Brownian motion and $\gamma_{n}^{2}(s)=\frac{\partial\nu_{n}^{2}(s)}{\partial s}$. We interpret $\frac{G_{n}^{\star}(T)}{G_{n}^{\star}(0)}$ as a change of measure on $\mathcal{F}_{T}$. Under the new induced measure $\mathbb{Q}^{G_{n}}$, $\tilde{W}_{G_{n}}(T)=W_{G_{n}}(T)-\int_{0}^{T}\gamma_{n}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$ is a Brownian motion. Then $G_{n}^{\star}(T)$ under $\mathbb{Q}^{G_{n}}$ has the same law as $G_{n}^{\star}(T)e^{\nu_{n}^{2}(T)}$ under $\mathbb{Q}$. Thus,
(24) $\mathbb{E}\left[{G_{n}}^{\star}(T)h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]$ $=\mathbb{E}^{G_{n}}\left[\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu_{np}^{2}(T)-\nu_{n}^{2}(T)}-\kappa^{\star}\right)^{+}\right]\,,$
where
(25) $\nu_{np}^{2}(T)=-\sum_{i=1}^{m_{p}-1}\sum_{j=m_{p}}^{m}a_{i}a_{j}\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
We have
(26) $\nu^{2}=\nu_{n}^{2}+\nu_{p}^{2}-2\nu_{np}^{2}\,.$
We know that ${G_{p}}^{\star}/G_{n}^{\star}$ has variance $\nu^{2}$. Furthermore, $G_{n}^{\star}$ and $G_{p}^{\star}$ are martingales of unit expectation under $\mathbb{Q}$, thus
(27) $\mathbb{E}\left[\frac{{G_{p}}^{\star}}{{G_{n}}^{\star}}(T)\right]=e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}}\,.$
Then using Equations 26 and 24, we obtain the result.
Identity 2.
(28) $\mathbb{E}\left[B(0,T)G_{p}^{\star}(T)h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\mathrm{Black}\left(e^{\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We perform a change of measure $\frac{G_{p}^{\star}(T)}{G_{p}^{\star}(0)}$. Under this new measure $\mathbb{Q}^{G_{p}}$, $\tilde{W}_{G_{p}}(T)=W_{G_{p}}(T)-\int_{0}^{T}\gamma_{p}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$, with $\gamma_{p}^{2}(s)=\frac{\partial\nu_{p}^{2}(s)}{\partial s}$ is a Brownian motion. Then $\frac{G_{p}}{G_{n}}^{\star}(T)$ under $\mathbb{Q}^{G_{p}}$ has the same law as $\frac{G_{p}}{G_{n}}^{\star}(T)e^{\nu_{p}^{2}-\nu_{pn}(T)}$ under $\mathbb{Q}$. Thus,
(29) $\mathbb{E}\left[G_{p}^{\star}(T)h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]$ $=\mathbb{E}^{G_{p}}\left[h\left(\frac{{G_{p}}^{\star}}{G_{n}^{\star}}(T)e^{\nu_{p}^{2}(T)-\nu_{np}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
Using Equation 26, we find that the right hand side corresponds to a vanilla option price under the Black model with forward $e^{\nu^{2}(T)}$, strike $\kappa^{\star}$ and total variance $\nu^{2}(T)$.
Identity 3.
(30) $\mathbb{E}\left[S_{i}^{\star}(T)h\left(\frac{G_{p}^{\star}}{G_{n}^{\star}}-\kappa^{\star}\right)\right]=\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\,.$
with
(31) $\bar{v}_{i}=\sum_{l=1}^{n}a_{l}v_{i,l}\,,$ $\quad v_{i,l}=\rho_{i,l}\int_{0}^{T}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
In particular, for an Asian, we have $v_{i,l}=\int_{0}^{t_{i}\wedge t_{j}}\sigma^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$. For baskets, the volatilities are taken to be constant equal to each underlying implied volatility to maturity in practice and we have $v_{i,l}=\rho_{i,l}\sqrt{v_{i}v_{l}}$.
Proof.
We use a change of measure $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$ and define
$\tilde{W}_{i}(t)=W_{i}(t)-\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$
to obtain
(32) $\mathbb{E}\left[S_{i}^{\star}(T)h\left(\frac{G_{p}^{\star}}{G_{n}^{\star}}-\kappa^{\star}\right)\right]$ $=\mathbb{E}^{i}\left[h\left(\frac{{G_{p}}^{\star}}{G_{n}^{\star}}e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-\kappa^{\star}\right)\right]\,.$
From Equation 27, we find that the right hand side corresponds to a vanilla option price under the Black model with forward $e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}+\bar{v}_{i}}$, strike $\kappa^{\star}$ and total variance $\nu^{2}(T)$.
Identity 4.
(33) $\mathbb{E}\left[B(0,T)G_{p}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=e^{\nu^{2}(T)}\mathrm{Black}\left(e^{2\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
Let $G^{\star}(T)=\frac{1}{\mathbb{E}\left[\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}\right]}\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}$, we perform a change of measure $\frac{G^{\star}(T)}{G^{\star}(0)}$. Under this new measure $\mathbb{Q}^{G}$, $\tilde{W}_{G}(T)=W_{G}(T)-\int_{0}^{T}\gamma(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$, with $\gamma^{2}(s)=\frac{\partial\nu^{2}(s)}{\partial s}$ is a Brownian motion. Then $\frac{G_{p}}{G_{n}}^{\star}(T)$ under $\mathbb{Q}^{G_{p}}$ has the same law as $\frac{G_{p}}{G_{n}}^{\star}(T)e^{\nu^{2}(T)}$ under $\mathbb{Q}$. Thus,
$\mathbb{E}\left[G_{p}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}\mathbb{E}^{G}\left[G_{p}^{\star}(T)e^{\nu_{p}^{2}-\nu_{pn}^{2}(T)}h\left(\frac{{G_{p}}^{\star}}{G_{n}^{\star}}(T)e^{\nu^{2}(T)}-\kappa^{\star}\right)\right]\,.$
Using Equation 26 and Identity 2 we find that the right hand side corresponds to a vanilla option price under the Black model with forward $e^{2\nu^{2}(T)}$, strike $\kappa^{\star}$ and total variance $\nu^{2}(T)$.
Identity 5.
(34) $\mathbb{E}\left[B(0,T)S_{i}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\mathrm{Black}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We start with a change of measure $\frac{G^{\star}(T)}{G^{\star}(0)}$:
$\mathbb{E}\left[B(0,T)S_{i}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}\mathbb{E}^{G}\left[B(0,T)S_{i}^{\star}(T)e^{\bar{v}_{i}}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu^{2}T}-\kappa^{\star}\right)\right]$
We then apply the change of measure $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$:
$\mathbb{E}\left[B(0,T)S_{i}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}+\bar{v}_{i}(T)}\mathbb{E}^{iG}\left[B(0,T)h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu^{2}T+\bar{v}_{i}}-\kappa^{\star}\right)\right]\,.$
Using Equations 26 and 27, we find that the right hand side corresponds to a vanilla option price under the Black model with forward $e^{\frac{-\nu_{p}^{2}+\nu_{n}^{2}+3\nu^{2}(T)}{2}+\bar{v}_{i}(T)}$, strike $\kappa^{\star}$ and total variance $\nu^{2}(T)$.
Identity 6.
(35) $\mathbb{E}\left[B(0,T)\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}(T)-2\nu_{np}^{2}+2\nu_{n}^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
with
(36) $\bar{v}_{i}^{-}(T)=\sum_{l=1}^{m_{p}-1}a_{l}\rho_{li}\int_{0}^{T}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,,$
and $v_{i,j}$ defined in Equation 31.
Proof.
We have $\mathbb{E}\left[\frac{1}{G_{n}^{\star}(T)}\right]=e^{\nu_{n}^{2}(T)}$. Let $G_{n}^{-1\star\star}=\frac{e^{-\nu_{n}^{2}(T)}}{G_{n}^{\star}(T)}$, it is a martingale of unit expectation under $\mathbb{Q}$ and may be thus expressed as
[…]
in probability, with $W_{G_{n}^{-1}}$ being a Brownian motion and $\gamma_{n}^{2}(s)=\frac{\partial\nu_{n}^{2}(s)}{\partial s}$. We interpret $\frac{G_{n}^{-1\star\star}(T)}{G_{n}^{-1\star\star}(0)}$ as a change of measure on $\mathcal{F}_{T}$. Under the new induced measure $\mathbb{Q}^{G_{n}^{-1}}$, $\tilde{W}_{G_{n}^{-1}}(T)=W_{G_{n}^{-1}}(T)-\int_{0}^{T}\gamma_{n}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$ is a Brownian motion. Then
$\mathbb{E}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}(T)}\mathbb{E}^{G_{n}^{-1}}\left[S_{i}^{\star}(T)S_{j}^{\star}(T)e^{\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
We then apply the change of measure $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$:
$\mathbb{E}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}(T)}\mathbb{E}^{iG_{n}^{-1}}\left[S_{j}^{\star}(T)e^{{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{\bar{v}_{i}-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
And finally a change of measure $\frac{S_{j}^{\star}(T)}{S_{j}^{\star}(0)}$:
$\mathbb{E}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)}{G_{n}^{\star}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\mathbb{E}^{jiG_{n}^{-1}}\left[h\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{\bar{v}_{i}+\bar{v}_{j}-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
Using Equations 26 and 27, we find that the right hand side corresponds to a vanilla option price under the Black model with forward $e^{\bar{v}_{i}+\bar{v}_{j}(T)-2\nu_{np}^{2}+2\nu_{n}^{2}(T)}$.
Identity 7.
(37) $\mathbb{E}\left[B(0,T)G_{p}^{\star}(T)\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}\right)^{2}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=e^{3\nu^{2}(T)}\mathrm{Black}\left(e^{3\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We start with the change of measure $\frac{G^{\star}(T)}{G^{\star}(0)}$
$\mathbb{E}\left[B(0,T)G_{p}^{\star}(T)\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}\right)^{2}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}\mathbb{E}\left[B(0,T)G_{p}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu_{p}^{2}-\nu_{pn}^{2}+\nu^{2}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu^{2}(T)}-\kappa^{\star}\right)\right]\,.$
We then apply Identity 4 with the additional overall multiplicative factor
$e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}e^{\nu_{p}^{2}-\nu_{pn}^{2}+\nu^{2}(T)}=e^{2\nu^{2}(T)}\,,$
and forward multiplier $e^{\nu^{2}(T)}$.
Identity 8.
(38) $\mathbb{E}\left[B(0,T)S_{i}^{\star}(T)\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}\right)^{2}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu^{2}+2\nu_{n}^{2}-2\nu_{pn}^{2}+2\bar{v}_{i}}\mathrm{Black}\left(e^{2\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We start with a change of measure $\frac{G^{\star}(T)}{G^{\star}(0)}$:
$\mathbb{E}\left[B(0,T)S_{i}^{\star}(T)\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}\right)^{2}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}\mathbb{E}^{G}\left[B(0,T)S_{i}^{\star}(T)\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\bar{v}_{i}+\nu^{2}(T)}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu^{2}T}-\kappa^{\star}\right)\right]$
We then apply Identity 5 with global multiplier
$e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}e^{\bar{v}_{i}+\nu^{2}(T)}$
and forward multiplier $e^{\nu^{2}T}$.
Identity 9.
(39) $\mathbb{E}\left[B(0,T)\frac{S_{i}^{\star}S_{j}^{\star}G_{p}^{\star}(T)}{{G_{n}^{\star}}^{2}}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{3\nu_{n}^{2}-2\nu_{pn}^{2}+\bar{v}_{i}+\bar{v}_{j}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}(T)+\nu^{2}-2\nu_{pn}^{2}+2\nu_{n}^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We start with a change of measure $\frac{G^{\star}(T)}{G^{\star}(0)}$:
$\mathbb{E}\left[B(0,T)\frac{S_{i}^{\star}S_{j}^{\star}G_{p}^{\star}(T)}{{G_{n}^{\star}}^{2}}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}\mathbb{E}^{G}\left[B(0,T)\frac{S_{i}^{\star}(T)S_{j}^{\star}}{G_{n}^{\star}(T)}e^{\bar{v}_{i}+\bar{v}_{j}+\nu_{n}^{2}(T)-\nu_{pn}^{2}}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}e^{\nu^{2}T}-\kappa^{\star}\right)\right]$
We then apply Identity 6 with global multiplier
$e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}(T)}{2}}e^{\bar{v}_{i}+\bar{v}_{j}+\nu_{n}^{2}(T)-\nu_{pn}^{2}}$
and forward multiplier $e^{\nu^{2}T}$.
Identity 10.
(40) $\mathbb{E}\left[B(0,T)\frac{S_{i}^{\star}S_{j}^{\star}S_{k}^{\star}(T)}{{G_{n}^{\star}}^{2}}h\left(\frac{G_{p}^{\star}(T)}{G_{n}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{{v}_{i,j}+{v}_{k,i}+{v}_{k,j}+2\bar{v}_{i}^{-}+2\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+3\nu_{n}^{2}(T)}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{k}-3\nu_{np}^{2}+3\nu_{n}^{2}(T)},\kappa^{\star},\nu^{2},T\right)\,.$
Proof.
We start with a change of measure $\frac{G_{n}^{-1\star\star}(T)}{G_{n}^{-1\star\star}(0)}$:
$\mathbb{E}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)S_{k}^{\star}(T)}{{G_{n}^{\star}(T)}^{2}}h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}(T)}\mathbb{E}^{G_{n}^{-1}}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)S_{k}^{\star}(T)}{{G_{n}^{\star}(T)}}e^{\bar{v}_{i}^{-}+\bar{v}_{j}^{-}+\bar{v}_{k}^{-}(T)+\nu_{n}^{2}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
We then perform a change of measure $\frac{S_{k}^{\star}(T)}{S_{k}^{\star}(0)}$:
$\mathbb{E}\left[\frac{S_{i}^{\star}(T)S_{j}^{\star}(T)S_{k}^{\star}(T)}{{G_{n}^{\star}(T)}^{2}}h\left(\frac{{G_{p}}^{\star}(T)}{{G_{n}}^{\star}(T)}-\kappa^{\star}\right)\right]=\\ e^{\nu_{n}^{2}(T)}\mathbb{E}^{kG_{n}^{-1}}\left[\frac{S_{i}^{\star}S_{j}^{\star}(T)}{{G_{n}^{\star}(T)}}e^{{v}_{k,i}+{v}_{k,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+\nu_{n}^{2}(T)}h\left(\frac{{G_{p}}^{\star}(T)}{G_{n}^{\star}(T)}e^{\bar{v}_{k}-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}-\kappa^{\star}\right)\right]\,.$
And we apply Identity 6 with global multiplier
$e^{{v}_{k,i}+{v}_{k,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+2\nu_{n}^{2}(T)}$
and forward multiplier $e^{\bar{v}_{k}-\nu_{np}^{2}(T)+\nu_{n}^{2}(T)}$.
Appendix B Derivatives of the Black-76 formula towards the strike
The following formulae are used to compute the expansions in practice:
(41) $\frac{\partial}{\partial K}\mathrm{Black}(F,K,v,T)$ $=-\eta B(0,T)\Phi(\eta d_{2})\,,$
(42) $\frac{\partial^{2}}{\partial K^{2}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K\sqrt{v}}\phi(d_{2})\,,$
(43) $\frac{\partial^{3}}{\partial K^{3}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K^{2}\sqrt{v}}\left(\frac{\phi(d_{2})}{\sqrt{v}}-1\right)\,,$
where $\Phi$, $\phi$ are the cumulative normal distribution function and the normal probability density function.